# Gradient descent avoids strict saddles with a simple line-search method too

### v1

@@ Abstract
It is known that gradient descent (GD) on a $C^{2}$ cost function generically avoids strict saddle points when using a small, constant step size. However, no such guarantee existed for GD with a line-search method. We provide one for a modified version of the standard Armijo backtracking method with generic, arbitrarily large initial step size. In contrast to previous works, our analysis does not require a globally Lipschitz gradient.
@@ 1 Introduction
Consider minimizing $f\colon{\mathbb{R}^{n}}\to{\mathbb{R}}$ by iterating $x_{t+1}=x_{t}-\alpha_{t}\nabla f(x_{t})$. A standard backtracking line-search sets an initial step size $\bar{\alpha}>0$, a decay parameter $\tau\in(0,1)$ and a tolerance parameter $r\in(0,1)$, then selects the step size $\alpha_{t}$ as the largest value in the geometric sequence $\bar{\alpha},\tau\bar{\alpha},\tau^{2}\bar{\alpha},\dots$ that satisfies the sufficient decrease condition $f(x_{t})-f(x_{t+1})\geq r\alpha_{t}\|\nabla f(x_{t})\|^{2}$—see (Nocedal and Wright, 2006, Alg. 3.1).
@@
Alternatively, one can express the update rule as $x_{t+1}=g_{i(x_{t})}(x_{t})$, where each $g_{i}(x)=x-\tau^{i}\bar{\alpha}\nabla f(x)$ is smooth, and the index $i(x_{t})$ selects the smallest exponent $i$ satisfying the sufficient decrease condition. This viewpoint retains smoothness of each candidate iteration map, but introduces a new complication: the selection $i(x_{t})$ depends on the current iterate $x_{t}$ which itself depends on $x_{0}$. Consequently, the algorithm does not apply a pre-specified sequence of iteration maps (common to all $x_{0}$). As such, classical analyses of non-adaptive dynamical systems do not apply.
In both formulations, the essential difficulty is the same: the iteration lacks the regularity or structure assumed by the CSMT and related tools. These difficulties might be resolved if, for example, the selected step-sizes eventually become constant or periodic, but this does not seem to be the case with the standard line-search described above.
In order to make some progress, we focus on the second viewpoint with a slightly modified procedure (Algorithm 1): rather than restarting the line-search from a fixed step size $\bar{\alpha}$ at each iteration, we use the previous step size $\alpha_{t-1}$ as the first candidate at iteration $t$. This modification has a significant effect: it ensures that if the algorithm converges, then the step size eventually stabilizes. Once the step size is constant, the update rule reduces to a smooth, fixed-step-size map. Exploiting this perspective, our main Euclidean result is as follows.
@@
Details including a generalization to Riemannian manifolds are given in Section 4. To get there, we first provide general results about avoidance of unstable fixed points in dynamical systems (Section 2), and improved guarantees for GD with constant step size (Section 3). Afterwards, we provide additional results for the Riemannian case with constant (not necessarily generic) step size (Sections 5 and 6).
@@
At every iteration, Algorithm 1 computes the next point as $x_{t+1}=g_{i_{t}}(x_{t})$ with some map $g_{i_{t}}$ selected from the collection of maps $g_{i}(x)=x-\tau^{i}\bar{\alpha}\nabla f(x)$ where $i\geq 0$ is an integer. If the sequence of points $x_{0},x_{1},x_{2},\ldots$ converges, then the step sizes decrease (that is, the indices increase, $i_{0}\leq i_{1}\leq i_{2}\leq\cdots$) and (we shall argue) eventually stabilize.
Thus, the algorithm has two phases. In the early phase, before the step size stabilizes, the algorithm selects a sequence of iteration maps that may vary from step to step. This sequence is not predetermined: it depends on $x_{0}$ in a discontinuous fashion. In the later phase, once the accepted step size becomes constant, the algorithm reduces to gradient descent with a fixed step size. In that regime, the tools from smooth dynamical systems such as the CSMT apply.
Accordingly, the analysis has two parts. First, we assert local control using the CSMT. Explicitly, we show that there exists a set $W$ of measure zero such that, if the algorithm converges to a strict saddle, then the iterates eventually all belong to $W$. Second, we globalize the result by noting that if some iterate $x_{t}$ belongs to $W$, then $x_{t-1}$ belongs to the preimage of $W$ by the iteration map that was chosen at step $t-1$. If that map has the property that preimages of measure-zero sets have measure zero—called the Luzin $N^{-1}$ property (Definition 2.4)—then we see that $x_{t-1}$ also belongs to a measure zero set. We do not know which iteration map was chosen at step $t-1$, but there are only countably many options, which is good enough. Iterating this reasoning, we find that $x_{0}$ itself belongs to a measure-zero set, as desired.
@@
The classical proof assumes that $\nabla f$ is $L$-Lipschitz continuous, and that the constant step size $\alpha$ satisfies $\alpha<1/L$. However, the main purpose of a line-search method is to avoid such impractical restrictions on the step size. Thus, the step sizes we entertain may be far larger than previously allowed, and that is true of the eventual (stabilized) step size as well.
Large step sizes are incompatible with the usual version of the CSMT used in this literature (Shub, 1987, Thm. III.7). To resolve this, we call upon a better (and just as classical) version of the CSMT (Hirsch et al., 1977, Thm. 5.1). For our purposes, they differ mainly in that the latter does not require the differential of the iteration map to be invertible anywhere. As discussed in Section 2, the Luzin $N^{-1}$ property only requires that differential to be invertible almost everywhere, without requirements at any specific points (not even at the saddle points themselves). This yields the following general statement, proved in Section 2. Compare with (Lee et al., 2019, Thm. 2) which requires $\mathrm{D}g(x)$ to be invertible for all $x$.
@@
By this reasoning, we can secure both local control and the local-to-global extension upon ensuring that the individual iteration maps each have a differential that is invertible almost everywhere. We do so using a Fubini–Tonelli-type argument. This requires the initial step size $\bar{\alpha}$ to be generic (righteously so, see Remark 4.2), but it can be arbitrarily large. In particular, we no longer need to assume the gradient of $f$ to be Lipschitz continuous.
@@
On a real-analytic manifold, if the Riemannian metric is also real-analytic (which is common though not required by Theorem 1.4), then the exponential map (see Section 5) is real-analytic everywhere on its domain. This is because the exponential map emerges from solving real-analytic differential equations, which have real-analytic solutions by the Cauchy–Kowalewski theorem (Krantz and Parks, 2002, §2.8).
@@
The base result by Lee et al. (2019) regards GD with constant step-size $\alpha<1/L$, where $\nabla f$ is $L$-Lipschitz continuous. They also obtain results for other discrete optimization algorithms including coordinate descent, block coordinate descent, mirror descent and Riemannian gradient descent with metric projection retraction. Li et al. (2019) apply similar ideas to study alternating minimization and proximal alternating minimization.
@@
Panageas et al. (2019) show that GD (among others) with step sizes $\alpha_{t}$ decaying as $\Omega(1/t)$ avoids isolated strict saddle points. Schaeffer and McCalla (2020, Cor. 3.8) consider GD with a pre-set sequence of piecewise constant step sizes. More specifically, the step size $\alpha_{t}$ at time $t$ is a piecewise constant function of $t$ and does not depend on the initial point. Additionally, the result requires that the gradient of the cost is globally $L$-Lipschitz continuous and that $\alpha_{t}L\in(0,2)$ for all times $t\geq 0$, which does not hold in practice for line-search methods. A similar result is presented by Shi and Iyengar (2020, Thm. 7.2) who, under a global Lipschitz Hessian condition and the assumption that all strict saddles are isolated, show that strict saddle points are avoided when using a pre-set sequence of step sizes that conform with the local Lipschitz constant of the gradient.
@@ 2 Unstable fixed point avoidance for dynamical systems
In the context of this paper, the strict saddles of a cost function $f$ would be unstable fixed points of an algorithm with iteration map $g$.
@@ 2.1 A general unstable fixed point avoidance theorem
for all $\textrm{for all }E\subseteq\mathcal{M},$ $\mu(E)=0\quad\implies\quad\mu(g^{-1}(E))=0.$ (Luzin $N^{-1}$)
@@
Notice that $f$ is smooth and satisfies the requirements above. Now, consider the map $g(x)=x-\nabla f(x)$. The origin is an unstable fixed point for $g$ since $g(0)=0$ and $\mathrm{D}g(0)$ has eigenvalues 00 and $2$. (The origin is also a strict saddle for $f$.) Moreover, for all $x\in{\mathbb{R}}^{2}$ with $\|x\|>2$, we have $g(x)=0$ so that $g$ does not avoid its unstable fixed points, and indeed it does not have the Luzin $N^{-1}$ property.
@@ 2.2 Proof of the general theorem
It is worth noting that the full CSMT provides additional information we won’t need, such as the regularity of $W_{\mathrm{cs}}^{\mathrm{loc}}$ near $x^{\ast}$. Moreover, it should be emphasized that the theorem is informative even for trajectories that converge to a point other than $x^{\ast}$ yet within the neighborhood $B$. This is particularly useful in cases where unstable fixed points are non-isolated.
@@ 2.3 Verifying the Luzin $N^{-1}$ property
For completeness, we provide a proof of Theorem 2.10 in Appendix A.1. The proof is adapted from (Ponomarev, 1987) and relies on Lindelöf’s lemma and Sard’s theorem as the key technical ingredients. Notice that Theorem 1.3 is a corollary of Theorems 2.5 and 2.10.
@@ 3.1 The Euclidean case
For a cost function $f\colon{\mathbb{R}^{n}}\to{\mathbb{R}}$, the iteration map of gradient descent (GD) with constant step size $\alpha$ is $g_{\alpha}(x)=x-\alpha\nabla f(x)$. Theorem 2.5 applied to $g_{\alpha}$ provides the following corollary.
Corollary 3.1.
@@
We make a few remarks, then dedicate the rest of the section to the proof.
@@
Notice that Corollary 3.1 applies even if the function value is not monotonically decreasing along the trajectory. In contrast, previous works assume that the step size satisfies $\alpha<1/L$ or $\alpha<2/L$, where $L$ is the Lipschitz constant of $\nabla f$.
@@
Corollary 3.1 holds for almost all positive $\alpha$, but indeed not for all. Consider again the function $f$ from Remark 2.6. If $\alpha=1$, then for all points $x\in{\mathbb{R}}^{2}$ with $\|x\|>2$, we have $g(x)=0$. In this case, the iteration map $g$ maps a set of positive measure directly to a strict saddle, showing that GD with step size $\alpha=1$ indeed does not avoid the strict saddles of $f$. Corollary 3.1 implies that the issue can be resolved by perturbing $\alpha$.
For gradient descent, it is straightforward to see that strict saddles are unstable, regardless of the step size used. Indeed, the differential of the iteration map $g_{\alpha}$ is $\mathrm{D}g_{\alpha}(x)=I_{n}-\alpha\nabla^{2}f(x)$ for all $x$. At a strict saddle, $\nabla^{2}f(x)$ has at least one negative eigenvalue, so for any step size $\alpha>0$, $\mathrm{D}g_{\alpha}(x)$ has at least one eigenvalue larger than $1$.
It remains to verify the Luzin $N^{-1}$ property. Once this is secured, Corollary 3.1 follows from Theorem 2.5. In the following, we demonstrate that, in the Euclidean case, the Luzin $N^{-1}$ property holds for almost any step size, with no additional assumptions.
@@
For completeness, we include a proof in Appendix A.2. With this in hand, we now argue that the iteration map of GD satisfies the Luzin $N^{-1}$ property for almost any step size.

### v2

@@ Abstract
It is known that gradient descent (GD) on a $C^{2}$ cost function generically avoids strict saddle points when using a small, constant step size. However, no such guarantee existed for GD with a line-search method. We provide one for a modified version of the standard Armijo backtracking method with generic, arbitrarily large initial step size. The proof underlines the double role of the Luzin $N^{-1}$ property for the iteration maps, and allows to forgo the habitual Lipschitz gradient assumption.
@@ 1 Introduction
Consider minimizing $f\colon{\mathbb{R}^{n}}\to{\mathbb{R}}$ by iterating $x_{t+1}=x_{t}-\alpha_{t}\nabla f(x_{t})$. A standard backtracking line-search sets an initial step size $\bar{\alpha}>0$, a decay parameter $\tau\in(0,1)$ and a tolerance parameter $r\in(0,1)$, then selects the step size $\alpha_{t}$ as the first value in the geometric sequence $\bar{\alpha},\tau\bar{\alpha},\tau^{2}\bar{\alpha},\dots$ that satisfies the sufficient decrease condition $f(x_{t})-f(x_{t+1})\geq r\alpha_{t}\|\nabla f(x_{t})\|^{2}$—see (Nocedal and Wright, 2006, Alg. 3.1).
@@
Alternatively, one can express the update rule as $x_{t+1}=g_{i(x_{t})}(x_{t})$, where each $g_{i}(x)=x-\tau^{i}\bar{\alpha}\nabla f(x)$ is smooth, and the index $i(x_{t})$ selects the smallest integer exponent $i$ satisfying the sufficient decrease condition. This viewpoint retains smoothness of each candidate iteration map, but introduces a new complication: the selection $i(x_{t})$ depends on the current iterate $x_{t}$ which itself depends on $x_{0}$. Consequently, the algorithm does not apply a pre-specified sequence of iteration maps (common to all $x_{0}$). As such, classical analyses of non-adaptive dynamical systems do not apply.
In both formulations, the essential difficulty is the same: the iteration lacks the regularity or structure assumed by the CSMT and related tools. These difficulties might be resolved if, for example, the selected step sizes eventually become constant or periodic, but this does not seem to be the case with the standard line-search described above.
In order to make some progress, we focus on the second viewpoint with a slightly modified procedure. In doing so, we aim to retain two fundamental aspects of line-search methods:
Function-class adaptivity.
When $\nabla f$ is globally $L$-Lipschitz, the method achieves almost the same (worst-case) performance as if $L$ were known and we set $\alpha_{t}=1/L$.
Local-geometry adaptivity.
The step size should automatically adjust to the local curvature of the function, without requiring any explicit estimates.
Guided by these principles, we propose the modified procedure in Algorithm 1. Given an arbitrary threshold $\varepsilon>0$, we perform a standard line-search whenever the gradient norm is at least $\varepsilon$. If the gradient norm falls below this threshold, rather than starting the line-search from the initial step size $\bar{\alpha}$ at the next iteration, we use the most recently selected step size $\alpha_{t}$ as the first candidate. This is the only departure from the standard Armijo backtracking line-search scheme, and it only affects the trajectory in the small-gradient regime. If the trajectory eventually escapes a saddle, the algorithm naturally returns to the standard line-search regime.
Input: $\tau,r\in(0,1),x_{0}\in\mathcal{M},\bar{\alpha}>0,\varepsilon>0,$ retraction $\mathrm{R}$, e.g., if $\mathcal{M}={\mathbb{R}}^{n}$, $\mathrm{R}_{x}(v)=x+v$
$\alpha_{0}\leftarrow\bar{\alpha}$
for $t=0,1,2,\dots$ do
   $\alpha_{t}\leftarrow\tau\alpha_{t}$
  end while
  $x_{t+1}\leftarrow\mathrm{R}_{x_{t}}(-\alpha_{t}\mathrm{grad}f(x_{t}))$
  if $\|\mathrm{grad}f(x_{t+1})\|\geq\varepsilon$ then
   $\alpha_{t+1}\leftarrow\bar{\alpha}$ $\triangleright$ Reset initial step size for standard line-search
  else
   $\alpha_{t+1}\leftarrow\alpha_{t}$ $\triangleright$ Maintain previously used step size instead
  end if
end for
As we shall see, the modification preserves both adaptivity notions above, while ensuring that if the algorithm converges, then the step size eventually stabilizes. Once the step size is constant, the update rule reduces to a smooth, fixed-step-size map. Exploiting this perspective, our main Euclidean result is as follows.
@@
Details (including a generalization to Riemannian manifolds) are given in Section 4. To get there, we first build up general results about avoidance of unstable fixed points in dynamical systems (Section 2), and improved guarantees for GD with constant step size (Section 3). Afterwards, we provide additional results for the Riemannian case with constant (not necessarily generic) step size (Sections 5 and 6).
@@
At every iteration, Algorithm 1 computes the next point as $x_{t+1}=g_{i_{t}}(x_{t})$ with some map $g_{i_{t}}$ selected from the collection of maps $g_{i}(x)=x-\tau^{i}\bar{\alpha}\nabla f(x)$ where $i\geq 0$ is an integer. If the sequence of points $x_{0},x_{1},x_{2},\ldots$ converges, then in a first phase, the step sizes can oscillate, but eventually they become non-increasing and (we shall argue) stabilize.
Thus, the algorithm has two phases. In the early phase, before the step size stabilizes, we run standard Armijo line-search, so the algorithm selects a sequence of iteration maps that may vary from step to step. This sequence is not predetermined: it depends on $x_{0}$ in a discontinuous fashion. In the later phase, once the accepted step size becomes constant, the algorithm reduces to gradient descent with a fixed step size. In that regime, the tools from smooth dynamical systems such as the CSMT apply.
Accordingly, the analysis has two parts. First, we assert local control using the CSMT. Explicitly, we show that there exists a set $W$ of measure zero such that, if the algorithm converges to a strict saddle, then the iterates eventually all belong to $W$. Second, we globalize the result by noting that if some iterate $x_{t}$ belongs to $W$, then $x_{t-1}$ belongs to the preimage of $W$ by the iteration map that was chosen at step t1t-1. If that map has the property that preimages of measure-zero sets have measure zero—called the Luzin $N^{-1}$ property (Definition 2.4)—then we see that $x_{t-1}$ also belongs to a measure zero set. We do not know which iteration map was chosen at step t1t-1, but there are only countably many options, which is good enough. Iterating this reasoning, we find that $x_{0}$ itself belongs to a measure-zero set, as desired.
@@
The classical proof assumes that $\nabla f$ is $L$-Lipschitz continuous, and that the constant step size $\alpha$ satisfies $\alpha<1/L$, or, in some extensions, $\alpha<2/L$ (see related work). However, the main purpose of a line-search method is to avoid such impractical restrictions on the step size. Thus, the step sizes we entertain may be far larger than previously allowed, and that is true of the eventual (stabilized) step size as well. Large step sizes are incompatible with the usual proof techniques for the following reasons:
  • The version of the CSMT used in this literature (Shub, 1987, Thm. III.7) requires the differential of the iteration map to be invertible at strict saddles.
  • The Luzin $N^{-1}$ property has so far only been established for GD with step sizes up to $2/L$, where $L$ is the Lipschitz constant of the gradient.
To resolve this, we call upon a better (and just as classical) version of the CSMT (Hirsch et al., 1977, Thm. 5.1) which lifts the aforementioned restriction. Then, as discussed in Section 2, if the differential of the iteration map is invertible almost everywhere, the Luzin $N^{-1}$ property holds without requirements at any specific points (not even at the saddle points themselves). This yields the following general statement about dynamical systems, proved in Section 2. Compare with (Lee et al., 2019, Thm. 2) which requires $\mathrm{D}g(x)$ to be invertible for all $x$.
@@
Then, using Fubini–Tonelli-type arguments, we show that the GD iteration map satisfies the Luzin $N^{-1}$ property for almost all choices of step size (Lemma 3.7). We later use this to show that for generic $\bar{\alpha}$, all maps considered by the line-search have the Luzin $N^{-1}$ property. In particular, unlike previous work, we no longer need to assume that $\nabla f$ is globally Lipschitz continuous.
@@
On a real-analytic manifold, if the Riemannian metric is also real analytic (which is common though not required by Theorem 1.4), then the exponential map (see Section 5) is real analytic everywhere on its domain. This is because it emerges from solving real-analytic differential equations, which have real-analytic solutions by the Cauchy–Kowalewski theorem (Krantz and Parks, 2002, §2.8).
@@
The base result by Lee et al. (2019) regards GD with constant step size $\alpha<1/L$, where $\nabla f$ is $L$-Lipschitz continuous. They also obtain results for other optimization algorithms including coordinate descent, block coordinate descent, mirror descent and Riemannian gradient descent with metric projection retraction. Li et al. (2019) apply similar ideas to study alternating minimization and proximal alternating minimization.
@@
Several works study saddle avoidance for GD with non-constant step sizes, assuming an a priori fixed sequence that is independent of the initial point $x_{0}$. Panageas et al. (2019) show that GD (among others) with step sizes $\alpha_{t}$ decaying as $\Omega(1/t)$ avoids isolated strict saddle points. Schaeffer and McCalla (2020, Cor. 3.8) consider GD with a pre-set sequence of piecewise constant step sizes. More specifically, the step size $\alpha_{t}$ at time $t$ is a piecewise constant function of $t$ with finitely many jumps. Additionally, the result requires that $\nabla f$ is globally $L$-Lipschitz continuous and that $\alpha_{t}L\in(0,2)$ for all times $t\geq 0$, which does not hold in practice for line-search methods.
Another result is presented by Shi and Iyengar (2020, Thm. 7.2), who show that GD avoids strict saddles when using a pre-set sequence of step sizes, assuming a globally Lipschitz Hessian and that all strict saddles are isolated. The step sizes must be fixed in advance, which prevents the method from adapting to the global problem class. In addition, even though they are predetermined, the step sizes must satisfy conditions that depend on the trajectory itself, specifically on the local Lipschitz constant of the gradient around the current iterate, which is not known a priori. Therefore, the method is not adaptive to function class or local geometry.
There have been other attempts at establishing saddle avoidance for GD with line-search-type methods. For example, Truong (2025) proposes a line search that requires an explicit approximation of the local Lipschitz constant of the gradient at each step, which is not easily implementable.
@@ 2 Unstable fixed point avoidance for dynamical systems
In the context of this paper, the strict saddles of a cost function $f$ would be unstable fixed points of an optimization algorithm for $f$ with iteration map $g$.
@@ 2.1 A general unstable fixed point avoidance theorem
$\textrm{for all }E\subseteq\mathcal{M},$ $\mu(E)=0\quad\implies\quad\mu(g^{-1}(E))=0.$ (Luzin $N^{-1}$)
@@
Notice that $f$ is smooth and satisfies the requirements above. Now, consider the map $g(x)=x-\nabla f(x)$. The origin is an unstable fixed point for $g$ since $g(0)=0$ and $\mathrm{D}g(0)$ has eigenvalues 0 and $2$. The origin is also a strict saddle for $f$. Moreover, for all $x\in{\mathbb{R}}^{2}$ with $\|x\|>2$, we have $g(x)=0$ so that $g$ does not avoid its unstable fixed points, and indeed it does not have the Luzin $N^{-1}$ property.
@@ 2.2 Proof of the general theorem
It is worth noting that the full CSMT provides additional information we do not need, such as the regularity of $W_{\mathrm{cs}}^{\mathrm{loc}}$ near $x^{\ast}$. Moreover, it should be emphasized that the theorem is informative even for trajectories that converge to a point other than $x^{\ast}$ yet within the neighborhood $B$. This is particularly useful in cases where unstable fixed points are non-isolated, as first noted by Panageas and Piliouras (2016).
@@ 2.3 Verifying the Luzin $N^{-1}$ property
For completeness, we provide a proof of Theorem 2.10 in Appendix A. The proof is adapted from (Ponomarev, 1987) and relies on Lindelöf’s lemma and Sard’s theorem as the key technical ingredients. Notice that Theorem 1.3 is a corollary of Theorems 2.5 and 2.10.
@@ 3.1 The Euclidean case
For a cost function $f\colon{\mathbb{R}^{n}}\to{\mathbb{R}}$, the iteration map of gradient descent (GD) with constant step size $\alpha$ is $g_{\alpha}(x)=x-\alpha\nabla f(x)$. The main result of this section is stated below.
Theorem 3.1.
@@
It is straightforward to see that strict saddles are unstable, regardless of the step size used. Indeed, the differential of the iteration map $g_{\alpha}$ is $\mathrm{D}g_{\alpha}(x)=I_{n}-\alpha\nabla^{2}f(x)$ for all $x$. At a strict saddle, $\nabla^{2}f(x)$ has at least one negative eigenvalue, so for any step size $\alpha>0$, $\mathrm{D}g_{\alpha}(x)$ has at least one eigenvalue larger than $1$.
To apply Theorem 2.5, we need to show that the iteration map satisfies the Luzin $N^{-1}$ property. We make a few remarks, then dedicate the rest of this section to showing that, in the Euclidean case, this is indeed true for almost any step size, with no additional assumptions.
@@
Notice that Theorem 3.1 applies even if the function value is not monotonically decreasing along the trajectory. In contrast, previous works assume that the step size satisfies $\alpha<1/L$ or $\alpha<2/L$, where $L$ is the Lipschitz constant of $\nabla f$, and hence $f(x_{t+1})\leq f(x_{t})$.
@@
Theorem 3.1 holds for almost all positive $\alpha$, but indeed not for all. Consider again the function $f$ from Remark 2.6. If $\alpha=1$, then for all points $x\in{\mathbb{R}}^{2}$ with $\|x\|>2$, we have $g(x)=0$. In this case, the iteration map $g$ maps a set of positive measure directly to a strict saddle, showing that GD with step size $\alpha=1$ indeed does not avoid the strict saddles of $f$. Theorem 3.1 implies that the issue can be resolved by perturbing $\alpha$.
@@
For completeness, we include a proof in Appendix B. With this in hand, we now argue that the iteration map of GD satisfies the Luzin $N^{-1}$ property for almost any step size.